A 1205 kg car is running a speed of 35 m/s along a level road. How large a constant retarding force is required to stop it in 15 s? 


1
Expert's answer
2022-03-18T09:08:36-0400

The Newton's second law says

"F=ma=m\frac{\Delta v}{\Delta t}"

"F=1205\times \frac{0-35}{15}=-2812\:\rm N"